OpenStudy (anonymous):

A single card is drawn from a standard deck of cards. Find the probabilities if the given information is known about the chosen card. A face card is a jack, queen, or king. P(jack/black)=

OpenStudy (anonymous):

There are two black jacks in a set of cards. There are 52 cards in a set.

OpenStudy (anonymous):

There is a basic rule in probability that, if all events are equally likely, then the probability of a given event is the number of ways that event can happen, divided by the total number of events.

OpenStudy (anonymous):

For example the probability of rolling a three in one dice roll is 1 / 6 because there's one way to roll a three, and 6 possible outcomes. 1 / 6.

OpenStudy (anonymous):

So, how many ways can you get a black jack? And how many outcomes are possible over all? Then divide the first, by the second.

OpenStudy (anonymous):

Let me explain in a esyer way! There are 2 red jacks out of 52 cards. 2/52 which would mean 1 out of 26 chance of getting a red jack. There would be 12 face cards out of 52 cards. That would be 3/13 cards. The queen is 1/3 for face cards. (1/3)(3/13) = 3/39 = 1/13. So, 3 out of 13 for finding a face card and 1 out of 13 for finding a queen. Hope this helps you ! :D